# Current Address: ]Kinlin School, Fanshawe College, 1001 Fanshawe College Blvd., London, Ontario, Canada

### v3

@@ Appendix C Additive-Type Correlated Models: Two Simple Examples
$\Theta(q)=\left\{\begin{array}[]{cr}0&~{}~{}\text{if}~{}q=0\\ {1\over{q}}&~{}~{}\text{if}~{}q\neq 0\end{array}\right..$ (143)
@@
+$\mathcal{M}^{(1)}_{+}(Y;q)$ $=$ $\left(\begin{array}[]{cc}0&~{}{q\over 2}\left(Y^{-{2q\over{q+1}}}-\Theta(q)\right)\\ {q\over 2}Y^{2}\Theta(q)&Y\end{array}\right)$ (146)
$=$ $\left(\begin{array}[]{cc}0&-{1\over 2}\\ 0&Y\end{array}\right)+q\left(\begin{array}[]{cc}0&~{}{1\over 2}Y^{-{2q\over{q+1}}}\\ {1\over 2}Y^{2}\Theta(q)&0\end{array}\right)$ (151)
@@
+$\mathcal{M}^{(2)}_{+}(Y;q)$ $=$ $\left(\begin{array}[]{cc}Y&~{}qY^{a}\left(Y^{-{q\over{q+1}}}+\Theta(q)\right)\\ qY^{b}\left(Y^{-{q\over{q+1}}}-\Theta(q)\right)&-Y\end{array}\right)$ (158)
$=$ $\left(\begin{array}[]{cc}Y&0\\ -Y^{b}&-Y\end{array}\right)+q\left(\begin{array}[]{cc}0&~{}Y^{a}\left(Y^{-{q\over{q+1}}}+\Theta(q)\right)\\ Y^{b-{q\over{q+1}}}&0\end{array}\right)$ (163)

### v4

@@ Appendix C Additive-Type Correlated Models: Two Simple Examples
$\Upsilon(q)=\left\{\begin{array}[]{cr}0&~~\text{if}~q=0\\ {1\over{q}}&~~\text{if}~q\neq 0\end{array}\right.$. (203)
@@
$\mathcal{M}^{(1)}_{+}(Y;q)$ $=$ $\left(\begin{array}[]{cc}0&~{q\over 2}\left(Y^{-{2q\over{q+1}}}-\Upsilon(q)\right)\\ {q\over 2}Y^{2}\Upsilon(q)&Y\end{array}\right)$ (206)
$=$ $\left(\begin{array}[]{cc}0&-{1\over 2}\\ 0&Y\end{array}\right)+q\left(\begin{array}[]{cc}0&~{1\over 2}Y^{-{2q\over{q+1}}}\\ {1\over 2}Y^{2}\Upsilon(q)&0\end{array}\right)$ (211)
@@
$\mathcal{M}^{(2)}_{+}(Y;q)$ $=$ $\left(\begin{array}[]{cc}Y&~qY^{a}\left(Y^{-{q\over{q+1}}}+\Upsilon(q)\right)\\ qY^{b}\left(Y^{-{q\over{q+1}}}-\Upsilon(q)\right)&-Y\end{array}\right)$ (218)
$=$ $\left(\begin{array}[]{cc}Y&0\\ -Y^{b}&-Y\end{array}\right)+q\left(\begin{array}[]{cc}0&~Y^{a}\left(Y^{-{q\over{q+1}}}+\Upsilon(q)\right)\\ Y^{b-{q\over{q+1}}}&0\end{array}\right)$ (223)